OpenStudy (pawanyadav):

Integrate √(4+x^2)/(x^6) dx

hartnn (hartnn):

Have you tried integration by parts? \(u = \sqrt{4+x^2 }, v = x^{-6}\)

OpenStudy (pawanyadav):

Yes, it' don't helping.

hartnn (hartnn):

hmm... x = 2 tan y ?

OpenStudy (pawanyadav):

OK checking it ....

OpenStudy (pawanyadav):

Then how we will solve for Denominator.....

hartnn (hartnn):

with that even i am stuck at cos^3 y/ sin^6 y with some constants outside

OpenStudy (pawanyadav):

Same here..

ganeshie8 (ganeshie8):

\[\dfrac{\cos^3y}{\sin^6y} = \cot^3y\csc^3y = \cot y (\csc^2y-1)\csc^3y\] let \(u =\csc y\)

hartnn (hartnn):

nice! would it be a challenge to bring x back?

ganeshie8 (ganeshie8):

sure it would be messy haha

OpenStudy (pawanyadav):

I can't reach my answer with it.

ganeshie8 (ganeshie8):

how ?

OpenStudy (pawanyadav):

OK I'm trying.....

OpenStudy (pawanyadav):

The cot y is confusing me....

OpenStudy (triciaal):